Suppose consumers will demand 60 units of a product when the price is $15
per unit and 36 units when the price is $19 each. Find the demand equation,
assuming that is linear. Find the price per unit when 42 units are demanded.

Answer by addingup(3677)

we need two ordered pairs (x1,y1) and (x2, y2)
we have that:
at $15 the demand is 60, and
at $19 the demand is 36
this gives us:
(15, 60) and (19, 36)
since m = (y2 - y1)/(x2 - x1), we have:
m = (36 - 60)/(19 - 15) = -24/4 = -6 (note that we have a negative slope, which is generally the case)
----------------------
y-Intercept: QD = -6P + b
(15, 60)
60 = (-6 *15) + b
60 = -90 + b
60 + 90 = b
b = 150 this is your y-intercept.
.
To make sure you got the right answer, try it with the other set of numbers:
(19, 36)
36 = (-6*19) + b
36 = -114 + b
b = 150
Sorry, I have to go. I hope you are able to do the rest by yourself.